# RealWebAssist: A Benchmark for Long-Horizon Web Assistance with Real-World Users

### v1

@@ Abstract
1 Introduction
As an integral part of people’s daily life, many of our everyday tasks are performed on the internet. With the tremendous advances in open-ended agents driven by large language models (LLMs) and vision-language models (VLMs), there has been increasing interest in engineering web-based agents that can assist humans with complex tasks on the web following humans’ instructions (Zheng et al., 2024a; Nakano et al., 2022). Recent works have demonstrated the promising performance of web-based agents on planning (Putta et al., 2024; Wang et al., 2024; Yao et al., 2023) and Graphical User Interface (GUI) grounding (Cheng et al., 2024; Wu et al., 2024b; Gou et al., 2024; Yang et al., 2024; Xu et al., 2024), across diverse websites, tasks, and GUI interfaces.
Despite these encouraging results, there have not been systematic studies on long-horizon web assistance with real-world users. Existing benchmarks (e.g., Zhou et al. (2023); Deng et al. (2024); Cheng et al. (2024); Yao et al. (2022); Jang et al. (2024)) typically focus on performing a task based on a single instruction. Additionally, the instructions in the current benchmarks were not collected from real-world users during natural web use sessions, lacking the realism of real user instructions. As a result, these benchmarks do not capture the full complexity of real-world users’ web behavior and instructions.
@@ 1 Introduction
The sequential instruction following tasks in our RealWebAssist benchmark reflect the natural human behavior on the web. First, real-world users may not initially know what they are looking for. Thus, they need to engage in information seeking on multiple web pages (e.g., step 1-2 in Figure 1), sometimes even across websites. Second, based on new information such as product reviews, users may change their minds (e.g., step 3). Third, users give simple instructions that are seemingly ambiguous out of the context but could be interpreted based on spatial and temporal context via pragmatic reasoning (Goodman & Frank, 2016; Fried et al., 2023). For instance, the third instruction in Figure 1 does not explicitly describe which product, but an intelligent assistant should be able to infer the true user intent and correctly select the product in the user’s mind. Lastly, in our benchmark, users can browse the websites and have the autonomy to make critical decisions (such as purchasing) on their own, which is complementary to existing benchmarks that focus on agents’ planning ability to fully complete the tasks without human involvement.
We systematically evaluated state-of-the-art models, including GUI grounding, VLMs, and large reasoning models (LRMs). Our experimental results indicated significant limitations of the existing models, including grounding, understanding user intents, reasoning about the spatial and temporal context, and adapting to user-specific routines that emerge from repeated interactions.
2 Related Works
@@
| SeeClick (Cheng et al., 2024) | ✗ | ✗ | ✓ | ✓ | ✗ | 1200+ |
| WebArena (Zhou et al., 2023) | ✗ | ✗ | ✗ | ✗ | ✗ | 812 |
| Mind2Web (Deng et al., 2024) | ✗ | ✗ | ✓ | ✗ | ✗ | 2000+ |
| WebLINX (Lù et al., 2024) | ✗ | ✓ | ✓ | ✗ | ✗ | 512 |
| VideoWebArena (Jang et al., 2024) | ✗ | ✗ | ✗ | ✗ | ✓ | 2021 |
| WebShop (Yao et al., 2022) | ✗ | ✗ | ✗ | ✗ | ✗ | 12087 |
@@
Web Agent Benchmarks. Existing web agent benchmarks primarily evaluate the performance of web agents on tasks with clearly defined, unambiguous instructions, often overlooking the complexities of real-world users’ behavior and their instructions to an AI assistant. On WebArena (Zhou et al., 2023), Mind2Web (Deng et al., 2024), and WebShop (Yao et al., 2022), an agent follows a single instruction to perform an isolated task. While they offer an evaluation of an agent’s planning capacity, they lack the evaluation of an agent’s ability to follow a long sequence of user instructions on long-horizon web tasks. There have also been GUI grounding benchmarks, such as SeeClick (Cheng et al., 2024), that focused on grounding simple instructions to clicking actions on webpages. These instructions only instruct web agents to click web elements rather than reaching a user goal (e.g., purchasing an item). WebLINX (Lù et al., 2024) features sequential instruction following. However, the instructions were generated by annotators who received detailed guidelines and extensive training, rather than by actual users. The resulting instructions do not capture the nuances and complexity of real-world user instructions that naturally emerge in interactions with an assistant. In contrast, our RealWebAssist benchmark consists of sequential instruction following tasks for assisting real-world users, providing a novel set of challenges necessary for long-horizon web assistance for real-world users. Table 1 summarizes key differences between RealWebAssist and prior benchmarks.
@@
Autonomous Web Agents. There have been many recent works on engineering autonomous web agents through retrieval augmented planning (Kim et al., 2024; Zhou et al., 2024; Wu et al., 2024a; He et al., 2024; Pan et al., 2024), finetuning (Hong et al., 2024; Gur et al., 2024; Deng et al., 2024; Pang et al., 2024; Zhang & Zhang, 2024), learning workflows (Zhang et al., 2023; Wang et al., 2024; Zheng et al., 2024b; Majumder et al., 2023; Cai et al., 2024), reinforcement learning (Liu et al., 2018; Shi et al., 2017; Nogueira & Cho, 2016; Humphreys et al., 2022), and combinations of these methods (Liu et al., 2023; Putta et al., 2024). These methods focus on planning capacity for a single task. However, there has not been much work on improving web agents’ ability to understand and follow real-world users’ sequential instructions on long-horizon tasks.
@@
GUI Grounding. One key ability for web agents in many assistance tasks is to ground instructions to clicking actions on a webpage. Recent works have explored VLM finetuning (e.g., Gou et al. (2024); Wu et al. (2024b); Yang et al. (2024)) as well as prompting pretrained VLMs with segmentations of web elements (e.g., Yang et al. (2023)) for enabling GUI grounding. These methods generate coordinates or bounding boxes on webpages to indicate where to click. They have only been trained on low-level instructions that clearly refer to web elements. It remains unclear if they can understand real-world user instructions that must be interpreted considering context or may refer to high-level goals.
3 RealWebAssist Benchmark
3.1 Problem Setup
@@
In RealWebAssist, a web agent has access to the current instruction, webpage (as a screenshot), and all the past interactions (previous instructions & screenshots of webpages). Since we are focusing on tasks on real-world websites, it is challenging to ensure safety as well as reproducibility in an interactive evaluation setting. Therefore, we adopt an offline evaluation setting following prior web-based agent benchmarks with real websites (Deng et al., 2024; Cheng et al., 2024). Specifically, for each instruction collected from the human experiment, the agent needs to identify the correct element to interact with by providing a coordinate or a bounding box to click on the webpage. A web agent’s action is considered correct if the clicking coordinate or the center of the bounding box they provide falls in the annotated correct regions on the webpage. If there are multiple steps corresponding to one instruction, we evaluate if the web agent’s actions for the same instruction are all correct.
Evaluation Metrics. We consider the following evaluation metrics: {addmargin}[-20pt]0pt
@@ 3.2 Dataset Construction
Setup. We recruited 10 participants (4 female, mean age = 20 years) from a US university campus to construct the dataset. All participants were native or fluent English speakers. Each participant completed a 40-minute real-world web assistance session. During the sessions, they were given a series of open-ended web tasks and asked to verbally instruct an experimenter, who operated the computer on their behalf, to complete the tasks. We use screen recordings and a high-quality USB microphone to record speech as raw data. The user study was approved by an institutional review board. Participants provided consent to have their voice recorded and included in this dataset.
@@
User Tasks. To increase the instruction diversity and realism, participants received general web-based tasks requiring active information seeking, sub-goal planning, and comparison among various options. We generate the task list by few-shot prompting GPT-4o with open-ended tasks, followed by manual filtering to ensure task quality and feasibility. These tasks have only general guidance, ensuring flexibility for personal decision-making. Example tasks include “Purchase an outfit for a formal event”, and “Plan a 5-day trip to Japan, booking both flights and hotels”. A full list of tasks can be found in the appendix A.3.1.
@@
Annotations. We manually labeled RealWebAssist data to ensure high-quality annotations. We first segmented the full recording into individual clips corresponding to each user’s instructions. In our benchmark, we disregard user speech unrelated to explicit instructions for the assistant, such as filler words or verbalized thought processes. For each instruction, we provide raw speech, speech transcript, webpage, and the correct regions to click (in the form of one or more bounding boxes). When there were multiple correct answers for the instructions (for instance, “can you close all the current tabs”), we annotated multiple bounding boxes as correct. When the experimenter made a mistake during the data collection sessions, we annotated the correct action intended by the user. If an instruction required multiple steps to complete, we set the instruction at each step as the same instruction. To generate the text instructions, we used an off-the-shelf recognition model, Whisper Large-V3 (Radford et al., 2023), to transcribe users’ speech and then manually fixed transcription errors. We provide more details in Appendix A.2.
Dataset Statistics. RealWebAssist consists of 1885 instructions spanning 2524 unique webpages. Each instruction corresponds to one or more webpages, depending on the number of steps needed to follow the instructions. There are 107 unique tasks across 66 different websites (a full list of websites and their task types is available in Appendix A.3.2). In addition to the annotated clips for individual instructions, we also plan to release the raw data, consisting of over 6 hours of video and speech. More dataset statistics are provided in Appendix A.3.
3.3 Key Challenges
@@
RealWebAssist features multiple challenges that could emerge in long-horizon web assistance with real-world users, many of which are not present in existing web agent benchmarks that only have clear, unambiguous, and non-sequential instructions, e.g., SeeClick (Cheng et al., 2024), WebArena (Zhou et al., 2023), and Mind2Web (Deng et al., 2024). Figure 3 illustrates the most common types of challenges, including spatial and temporal reasoning needed to understand ambiguous and context-dependent user instructions, planning for multiple steps of actions to reach the goal communicated by an instruction, and learning about user-specific routines. These key challenges provide a more realistic and holistic evaluation of a web agent’s reasoning, planning, and learning abilities to assist real-world users on long-horizon tasks.
@@
Temporal Reasoning. In our sequential instruction following tasks, users may instruct an assistant with the history as assumed temporal context. For example, to understand the intended meaning of “click the last item,” the assistant must memorize the items the user has viewed in the past. Figure 3B shows temporal reasoning based on different kinds of temporal context, ranging from short context between two consecutive webpages to long context with the same website to long context across websites. From the temporal context, the assistant needs to memorize crucial elements in the previous webpages, infer and track a user’s mind (e.g., change of mind about what to buy) based on the past instructions and webpages, and identify the earlier webpage the user refers to. Such temporal reasoning has not been evaluated in prior web agent benchmarks. However, it is very common in our benchmark due to the nature of human web browsing behavior as well as human instructions guided by pragmatics (Goodman & Frank, 2016).
@@
Routine. Since our benchmark allows a user to engage in repeated interactions with an assistant over multiple tasks, we observe that users may define routines understood by the assistant after repeated interactions. As shown in Figure 3D, the user initially gave detailed step-by-step instructions when selecting arrival and departure dates for a flight. In a subsequent task, however, the user simplified them into a single instruction when selecting dates for a hotel room. Such shorter instructions become possible after establishing a routine in the earlier task. Cognitive studies found that procedural abstraction, like these routines, naturally emerges in human cooperative communication through repeated interactions, allowing more efficient communication with partners (McCarthy et al., 2021). The emergence of such routines in our benchmark poses a novel challenge for web agents—learning user-specific procedural abstraction via repeated interactions to achieve human-like adaptive assistance. We hypothesize that this ability could enhance users’ perception of the AI assistant, as it understands human cooperative communication.
4 Experiments
4.1 Baselines
@@
We evaluated several types of models for web agents commonly evaluated in existing web agent benchmarks that have real-world websites (i.e., offline evaluation):
@@
GUI Grounding Models. GUI grounding models directly translate an instruction to an action on a webpage. There are two general types of grounding models. First, Set-of-Mark (SoM) (Yang et al., 2023) segments salient elements on a webpage using an off-the-shelf segmentation model (e.g., SAM (Kirillov et al., 2023) and Semantic-SAM (Li et al., 2023)) and prompts a VLM to select a segment mask as to identify the clicking area corresponding to the given instruction. Second, VLMs finetuned on datasets with paired instructions and annotated clicking coordinates or bounding boxes. We evaluated UGround-V1 (Gou et al., 2024), OS-Atlas (Wu et al., 2024b), and Aria-UI (Yang et al., 2024).
@@
VLM/LRM + Grounding. Grounding models are designed or trained to ground a simple instruction to a webpage and thus tend to lack reasoning or planning capabilities. To address this, we leveraged VLMs and LRMs to first translate real user instructions to more understandable ones for grounding models. In particular, a VLM or an LRM needs to reason about the true user intent implied by the instruction and the spatial & temporal context. For instructions that require multiple actions, it needs to generate a plan to complete the instructions. Finally it needs to generate a straightforward, clear instruction for the grounding model to produce the final action at each step. In the current benchmark, we evaluated state-of-the-art VLMs including GPT-4o (OpenAI, 2023), Gemini 2.0 Flash (Team et al., 2023) and Qwen-2.5 Instruct-72B (Qwen et al., 2025), as well as state-of-the-art LRMs including OpenAI o1 (Jaech et al., 2024), Gemini 2.0 Flash-Thinking (Team et al., 2023) and Claude 3.7 Sonnet (Anthropic, 2025). We paired all VLMs and LRMs with the best-performing grounding model (i.e., UGround-V1). For all VLMs and LRMs, we provide the past 10 steps for context, which we found to be a reasonable fixed context length in our preliminary study, balancing cost and informativeness. We also found that prompting models with screenshots of past webpages could incur a high cost. Therefore, we only prompt the models with the screenshot of the current webpage. For the history, we prompted GPT-4o to generate text-based action history based on consecutive screenshots and the instructions at each step. We then used this text-based history description for the evaluated VLMs and LRMs.
@@
Finetuning. To evaluate whether models can learn to better follow real-world user instructions using real-world user data, we finetuned the best-performing grounding model (UGround-V1) following (Zheng et al., 2024c) on 9 participants’ data and tested it on the held-out participants’ instructions. Specifically, we trained the grounding model to produce an action based on the past 10 steps of actions (in text), the current webpage screenshot, and the instruction. We enumerated different train/test splits and reported the averaged performance, either using the finetuned model alone or pairing it with the best VLM or LRM.
4.2 Results

### v2

@@ Abstract
Introduction
As an integral part of people’s daily life, many of our everyday tasks are performed on the internet. With the tremendous advances in open-ended agents driven by large reasoning models (LRMs) and vision-language models (VLMs), there has been increasing interest in engineering web-based agents that can assist humans with complex tasks on the web following humans’ instructions (zheng2024gpt; nakano2022webgptbrowserassistedquestionansweringhuman). Recent works have demonstrated the promising performance of web-based agents on planning (putta2024agent; wang2024agent; yao2023reactsynergizingreasoningacting) and Graphical User Interface (GUI) grounding (cheng2024seeclick; wu2024atlas; gou2024navigating; yang2024aria; xu2024aguvisunifiedpurevision), across diverse websites, tasks, and GUI interfaces.
Despite these encouraging results, there have not been systematic studies on long-horizon web assistance with real-world users. Existing benchmarks (e.g., (zhou2023webarena; deng2024mind2web; cheng2024seeclick; yao2022webshop; jang2024videowebarena)) typically focus on performing a task based on a single instruction. Additionally, the instructions in the current benchmarks were not collected from real users during natural web use sessions, lacking the realism of real user instructions. As a result, these benchmarks do not capture the full complexity of real users’ web behavior and instructions.
@@ Introduction
The sequential instruction following tasks in our RealWebAssist benchmark reflect the natural human behavior on the web. First, real-world users may not initially know what they are looking for. Thus, they need to engage in information seeking on multiple web pages (e.g., step 1-2 in Figure 1), sometimes even across websites. Second, based on new information such as product reviews, users may change their minds (e.g., step 3). Third, users give simple instructions that are seemingly ambiguous out of the context but could be interpreted based on spatial and temporal context via pragmatic reasoning (goodman2016pragmatic; fried2023pragmaticslanguagegroundingphenomena). For instance, the third instruction in Figure 1 does not explicitly describe which product, but an intelligent assistant should be able to infer the true user intent and correctly select the product in the user’s mind. Lastly, in our benchmark, users can browse the websites and have the autonomy to make critical decisions (such as purchasing) on their own, which is complementary to existing benchmarks that focus on agents’ planning ability to fully complete the tasks without human involvement.
We systematically evaluate state-of-the-art models, including GUI grounding, VLMs, and large reasoning models. Experimental results reveal that these models lack several key abilities, including grounding, understanding user intents, reasoning about spatial and temporal context, and adapting to user-specific routines.
Related Works
@@
| SreenSpot (cheng2024seeclick) | ✗ | ✗ | ✓ | ✓ | ✗ | 1200+ |
| WebArena (zhou2023webarena) | ✗ | ✗ | ✗ | ✗ | ✗ | 812 |
| Mind2Web (deng2024mind2web) | ✗ | ✗ | ✓ | ✗ | ✗ | 2000+ |
| WebLINX (lu2024weblinx) | ✗ | ✓ | ✓ | ✗ | ✗ | 512 |
| VideoWebArena (jang2024videowebarena) | ✗ | ✗ | ✗ | ✗ | ✓ | 2021 |
| WebShop (yao2022webshop) | ✗ | ✗ | ✗ | ✗ | ✗ | 12087 |
| BearCubs (song2025bearcubs) | ✗ | ✗ | ✓ | ✗ | ✗ | 111 |
@@
Web Agent Benchmarks. Existing web agent benchmarks primarily evaluate the performance of web agents on tasks with clearly defined, unambiguous instructions, often overlooking the complexities of real-world users’ behavior and their instructions to an AI assistant. On WebArena (zhou2023webarena), Mind2Web (deng2024mind2web), and WebShop (yao2022webshop), an agent follows a single instruction to perform an isolated task. While they offer an evaluation of an agent’s planning capacity, they lack the evaluation of an agent’s ability to follow a long sequence of user instructions on long-horizon web tasks. There have also been GUI grounding benchmarks, such as ScreenSpot (cheng2024seeclick), that focused on grounding simple instructions to clicking actions on webpages. These instructions only instruct web agents to click web elements rather than reaching a user goal (e.g., purchasing an item). WebLINX (lu2024weblinx) features sequential instruction following. However, the instructions were generated by annotators who received detailed guidelines and extensive training, rather than by actual users. The resulting instructions do not capture the nuances and complexity of real-world user instructions that naturally emerge in interactions with an assistant. In contrast, RealWebAssist consists of sequential instruction following tasks for assisting real-world users, providing a novel set of challenges necessary for long-horizon web assistance for real-world users. Table 1 summarizes key differences between RealWebAssist and prior benchmarks.
@@
Autonomous Web Agents. There have been many recent works on engineering autonomous web agents through retrieval augmented planning (kim2024rada; zhou2024languageagenttreesearch; wu2024oscopilotgeneralistcomputeragents; he2024webvoyagerbuildingendtoendweb; pan2024autonomousevaluationrefinementdigital), finetuning (hong2024cogagentvisuallanguagemodel; gur2024realworldwebagentplanninglong; deng2024mind2web; pang2024iterativereasoningpreferenceoptimization; zhang2024lookscreensmultimodalchainofaction), learning workflows (zhang2023appagentmultimodalagentssmartphone; wang2024agent; zheng2024synapsetrajectoryasexemplarpromptingmemory; majumder2023clincontinuallylearninglanguage; cai2024largelanguagemodelstool), reinforcement learning (liu2018reinforcement; shi2017world; nogueira2016end; humphreys2022data), and combinations of these methods (liu2023bolaa; putta2024agent). These works focus on planning for a single task. However, there has not been much work on understanding and following real-world users’ sequential instructions on long-horizon tasks.
@@
GUI Grounding. One key ability for web agents in many assistance tasks is to ground instructions to clicking actions on a webpage. Recent works have explored VLM finetuning (e.g., (gou2024navigating; wu2024atlas; yang2024aria; yang2025gta1; wu2025gui; qin2025ui; xu2025aguvisunifiedpurevision; yuan2025enhancingvisualgroundinggui)) as well as prompting pretrained VLMs with segmentations of web elements (e.g., (yang2023setofmark)) for enabling GUI grounding. These methods generate coordinates or bounding boxes on webpages to indicate where to click. They have only been trained on low-level instructions that clearly refer to web elements. It remains unclear if they can understand real-world user instructions that must be interpreted considering context or may refer to high-level goals.
RealWebAssist Benchmark
Problem Setup
@@
In RealWebAssist, a web agent has access to the current instruction, webpage (as a screenshot), and all the past interactions (previous instructions & screenshots of webpages). Since we are focusing on tasks on real-world websites, it is challenging to ensure safety and reproducibility in an interactive evaluation setting. Therefore, we adopt an offline evaluation setting following prior web-based agent benchmarks with real websites (deng2024mind2web; cheng2024seeclick). Specifically, for each instruction collected from the human experiment, the agent needs to identify the correct element to interact with by providing a coordinate or a bounding box to click on the webpage. As shown by figure 3, a web agent’s action is considered correct if the coordinate or the center of the bounding box they provide falls in the annotated correct regions on the webpage. If there are multiple steps corresponding to one instruction, we evaluate if the web agent’s actions for the same instruction are all correct.
Evaluation Metrics
We consider the following evaluation metrics:
@@
Setup. We recruited 10 participants (4 female, 6 male, mean age = 20 years) from a US university campus, none of whom had prior knowledge of the study’s purpose, to construct the dataset. All participants were native or fluent English speakers. Each participant completed a 40-minute real-world web assistance session in which they tackled a series of open-ended tasks designed to encourage diverse strategies. During each session, participants verbally instructed an experimenter, who operated the computer on their behalf, to complete the tasks. We captured screen recordings and used a high-quality USB microphone to record speech as raw data. The user study was approved by an institutional review board.
@@
User Tasks. To increase the instruction diversity and realism, participants received general web-based tasks requiring active information seeking, sub-goal planning, and comparison among various options. We generated the task list by few-shot prompting GPT-4o with open-ended tasks, followed by manual filtering and editing to ensure task quality and feasibility. These tasks provide only general guidance, ensuring flexibility for personal decision-making. Example tasks include “Purchase an outfit for a formal event” and “Plan a 5-day trip to Japan, booking both flights and hotels”. Each user finishes about 10 tasks.
Emergent User Behavior. In our realistic, open-ended settings, users exhibit rich behaviors that are not present in previous benchmarks. These include, but are not limited to, information seeking, researching and comparing different options, change of mind, and trial-and-error.
@@
Annotations. We manually labeled RealWebAssist data to ensure high-quality annotations. We first segmented the full recording into individual clips corresponding to each user’s instructions. In our benchmark, we disregard user speech unrelated to explicit instructions for the assistant, such as filler words or verbalized thought processes. For each instruction, we provide raw speech, speech transcript, webpage, and the correct regions to click (in the form of one or more bounding boxes). When there were multiple correct answers for the instructions (for instance, “can you close all the current tabs”), we annotated all correct regions with multiple bounding boxes. When the experimenter made a mistake during the data collection sessions, we annotated the correct action intended by the user. If an instruction required multiple steps to complete, we set the instruction at each step as the same instruction. To generate the text instructions, we used an off-the-shelf recognition model, Whisper Large-V3 (radford2023robust), to transcribe users’ speech and then manually fixed transcription errors. For all the instructions, we have three annotators verifying all of them, ensuring 100% agreement.
Dataset Statistics. RealWebAssist contains 1,885 user instructions across 107 tasks, 66 websites, and 2,524 screenshots. In addition to the benchmark, we also plan to release the raw data, consisting of over 6 hours of video & audio.
Key Challenges
@@
RealWebAssist features multiple challenges as illustrated in Figure 4, including spatial and temporal reasoning needed to understand ambiguous and context-dependent user instructions, planning for multiple steps of actions to reach the goal communicated by an instruction, and learning about user-specific routines. These key challenges provide a more realistic and holistic evaluation of a web agent’s reasoning, planning, and learning abilities to assist real-world users on long-horizon tasks. It is worth noting that many of these challenges, in particular, spatial reasoning, temporal reasoning, and routine understanding, are not present in existing web agent benchmarks. Unlike RealWebAssist, prior benchmarks, such as ScreenSpot (cheng2024seeclick), WebArena (zhou2023webarena), and Mind2Web (deng2024mind2web), only include clear, unambiguous, and non-sequential instructions.
@@
Temporal Reasoning. In our sequential instruction following tasks, users may instruct an assistant with the history as an assumed temporal context. For example, to understand the intended meaning of “click the last item,” the assistant must memorize the items the user has viewed in the past. Figure 4B shows temporal reasoning based on different kinds of temporal context, ranging from short context between two consecutive webpages to long context with the same website to long context across websites. From the temporal context, the assistant needs to memorize crucial elements in the previous webpages, infer and track a user’s mind (e.g., change of mind about what to buy) based on the past instructions and webpages, and identify the earlier webpage the user refers to. Such temporal reasoning has not been evaluated in prior web agent benchmarks. However, it is very common in our benchmark due to the nature of human web browsing behavior as well as human instructions guided by pragmatics (goodman2016pragmatic).
@@
Routine. Since our benchmark allows a user to engage in repeated interactions with an assistant over multiple tasks, we observe that users may define routines understood by the assistant after repeated interactions. As shown in Figure 4D, the user initially gave detailed step-by-step instructions when selecting arrival and departure dates for a flight. In a subsequent task, however, the user simplified them into a single instruction when selecting dates for a hotel room. Such shorter instructions become possible after establishing a routine in the earlier task. Cognitive studies found that procedural abstraction, like these routines, naturally emerges in human cooperative communication through repeated interactions, allowing more efficient communication with partners (mccarthy2021learning). The emergence of such routines in our benchmark poses a novel challenge for web agents—learning user-specific procedural abstraction via repeated interactions to achieve human-like adaptive assistance. We hypothesize that this ability could enhance users’ perception of the AI assistant, as it understands human cooperative communication.
Experiments
Baselines
@@
We evaluated several types of models for web agents commonly evaluated in existing web agent benchmarks that have real-world websites (i.e., offline evaluation). For all the experiments, we use the ground-truth captions for instructions.
@@
GUI Grounding Models. GUI grounding models directly translate an instruction to an action on a webpage. There are two general types of grounding models. First, Set-of-Mark (SoM) (yang2023setofmark) segments salient elements on a webpage using an off-the-shelf segmentation model (e.g., SAM (kirillov2023segment) and Semantic-SAM (li2023semantic)) and prompts a VLM to select a segment mask to identify the clicking area corresponding to the given instruction. Second, VLMs finetuned on datasets with paired instructions and annotated clicking coordinates or bounding boxes. We evaluated UGround-V1 (gou2024navigating), OS-Atlas (wu2024atlas), Aria-UI (yang2024aria), GTA-1 (yang2025gta1), GUI-Actor (wu2024oscopilotgeneralistcomputeragents), and UI-TARS (qin2025ui).
@@
VLM/LRM + Grounding. Grounding models are designed or trained to ground a simple instruction to a webpage and thus tend to lack reasoning or planning capabilities. To address this, we leveraged VLMs and LRMs to first translate real user instructions to more understandable ones for grounding models. In particular, a VLM or an LRM needs to reason about the true user intent implied by the instruction and the spatial & temporal context. For instructions that require multiple actions, it needs to generate a plan to complete the instructions. Finally, it needs to generate a straightforward, clear instruction for the grounding model to produce the final action at each step. We evaluated state-of-the-art VLMs (OpenAI2023GPT4TR; comanici2025gemini25pushingfrontier; qwen2025qwen25technicalreport), as well as state-of-the-art LRMs (jaech2024openai; comanici2025gemini25pushingfrontier; claude37sonnet2024). In the main results, we paired each VLM and LRM with the grounding model that achieved the highest step accuracy (GTA-1). For all VLMs and LRMs, we provide the past 10 steps for context, which we found to be a reasonable fixed context length in our preliminary study, balancing cost and informativeness. We also found that prompting models with screenshots of past webpages could incur a high cost. Therefore, we only prompt the models with the screenshot of the current webpage. For the history, we prompted GPT-4o to generate text-based action history based on consecutive screenshots and the instructions at each step. We then used this text-based history description for the evaluated VLMs and LRMs.
@@
Finetuning. To evaluate whether models can learn to better follow real-world user instructions with additional training, we finetuned the best-performing grounding model (GTA-1) following the model’s original group relative policy optimization (GRPO) training procedure (yang2025gta1) on 9 participants’ data and tested it on the held-out participants’ instructions. Specifically, we trained the grounding model to produce an action based on the past 10 steps of actions (in text), the current webpage screenshot, and the instruction. We enumerated different train/test splits and reported the averaged performance, either using the finetuned model alone or pairing it with the best VLM or LRM.
Results